Circular No. 7487

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

4U 0115+63 W. Coburn, R. Rothschild, and W. A. Heindl, University of California

at San Diego, report: “An outburst of the Be-transient x-ray pulsar 4U 0115+63 has been detected by the RXTE/ASM. The daily averaged ASM countrate began to increase on Aug. 27, which coincided with periastron of the source. In the first four days, the ASM countrate rose to > 7 counts/s. This is a faster rise than the 1999 Mar. outburst (IAUC 7116), which began several days after periastron, and a similar countrate was achieved 9 days after that. On 2000 Sept. 1.06 UT, the RXTE made a 1000-s pointed observation. Strong pulsations were detected at the 3.61-s pulse period, confirming that it is the pulsar currently in outburst. The 2–100-keV flux was (6.5 ± 0.3) × 10^{−9} erg cm^{−2} s^{−1}, and at a distance of 3500 pc had an x-ray luminosity of 9.6 × 10^{36} erg/s. Preliminary fitting of the joint PCA/HEXTE spectrum reveals two cyclotron absorption features at ∼ 15 and 23 keV, similar to what was observed the previous outburst (IAUC 7126). Because phase-averaged spectra are complex due to pulse-phase variability, these line energies are not well determined. Additional observations are encouraged. The RXTE/ASM real-time light curve is available

at http://heasarc.gsfc.nasa.gov/xte weather/.”

COMET 97P/METCALF-BREWINGTON G. V. Williams, Minor Planet Center, identified this comet with an

asteroidal LINEAR object, 1^{o}.1 from the prediction (MPC 31663). The nominal correction of ΔT = +3.5 days yields the ephemeris below but lacks literal meaning because of the 1993 passage < 0.11 AU from Jupiter.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Sept.1.26369  0^{h}27^{m}15.^{s}72  +6^{o}30′30′′.5  19.0
1.31736  0 27 14.72  +6 30 08.1  19.4
3.29703  0 26 39.81  +6 16 01.6  19.0
3.31054  0 26 39.61  +6 15 54.6  18.9
2000TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Sept. 3 0^{h}26.^{m}75 + 6^{o}18.^{′}3 2.106 3.034 152.^{o}0 9.^{o}0 19.6
13 0 23.04 + 4 57.7 2.024 3.001 163.1 5.6 19.4
23 0 18.26 + 3 24.0 1.968 2.969 174.6 1.8 19.1
Oct. 3 0 13.03 + 1 43.3 1.941 2.937 173.6 2.2 19.0
13 0 08.10 + 0 03.2 1.942 2.907 161.9 6.1 19.2
2000 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green